OpenStudy (anonymous):

Find the slope-intercept equation of the line with the following properties: Perpendicular to the line 푥 −8푦 = 8; containing the point (2,4).

OpenStudy (bahrom7893):

look who's here! what's up with the weird characters?

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

haha wow the symbol wasn't meant to be there! It is a copy/past problem. Anyway it should say x - 8y

OpenStudy (bahrom7893):

x-8y=8 -8y=8-x 8y=x-8 y=(1/8)x-1

OpenStudy (bahrom7893):

M = 1/8 perpendicular M = -8 (inverse negative)

OpenStudy (bahrom7893):

Your line is: y-4=-8(x-2) y-4=-8x+16 y=-8x+20

OpenStudy (anonymous):

so the final equation of y= -8+20 is the slope intercept form? I have not done this for a while and did not have time to read the chapter :'(

OpenStudy (bahrom7893):

yea y=-8x+20 is slope intercept form

OpenStudy (anonymous):

You're awesome! thanks :D

OpenStudy (bahrom7893):

np